# Spatial-temporal Memories Enhanced Graph Autoencoder for Anomaly Detection in Dynamic Graphs

### v1

@@ Abstract
Anomaly detection in dynamic graphs presents a significant challenge due to the temporal evolution of graph structures and attributes. The conventional approaches that tackle this problem typically employ an unsupervised learning framework, capturing normality patterns with exclusive normal data during training and identifying deviations as anomalies during testing. However, these methods face critical drawbacks: they either only depend on proxy tasks for general representation without directly pinpointing normal patterns, or they neglect to differentiate between spatial and temporal normality patterns, leading to diminished efficacy in anomaly detection. To address these challenges, we introduce a novel Spatial-Temporal memories-enhanced graph autoencoder (STRIPE). Initially, STRIPE employs Graph Neural Networks (GNNs) and gated temporal convolution layers to extract spatial features and temporal features, respectively. Then STRIPE incorporates separate spatial and temporal memory networks, which capture and store prototypes of normal patterns, thereby preserving the uniqueness of spatial and temporal normality. After that, through a mutual attention mechanism, these stored patterns are then retrieved and integrated with encoded graph embeddings. Finally, the integrated features are fed into the decoder to reconstruct the graph streams which serve as the proxy task for anomaly detection. This comprehensive approach not only minimizes reconstruction errors but also refines the model by emphasizing the compactness and distinctiveness of the embeddings in relation to the nearest memory prototypes. Through extensive testing, STRIPE has demonstrated a superior capability to discern anomalies by effectively leveraging the distinct spatial and temporal dynamics of dynamic graphs, significantly outperforming existing methodologies, with an average improvement of 15.39% on AUC values.
@@
Anomaly Detection, Dynamic Graphs, Graph Autoencoder, Memory Networks.
@@ I Introduction
Real-world networks are often modeled as dynamic graphs to reflect the changing nature of objects and their interactions [1, 2, 3, 4]. In addition to basic structural details and node attributes, dynamic graphs also contain rich temporal signals, e.g., the evolving patterns of the graph structure and node attributes [5, 6, 7]. This temporal aspect introduces an additional view for analyzing anomalies. For example, as shown in Fig. 1, the anomaly may not be evident when only considering the spatial information in the last graph snapshot since node 5 has similar structural information to node 3 in terms of node degrees and common neighbors. However, by observing the temporal changes in graph structures, the abnormality becomes distinctly noticeable. To avoid ambiguity, in this paper, we define the structures and attributive features within individual graph snapshots as spatial information, while characterizing the evolving changes and trends among different snapshots as temporal information. To model and integrate both the spatial and temporal signals of nodes and edges for anomaly detection, there has been a growing interest in the study of anomaly detection in dynamic graphs [8, 9].
Figure 1: A toy example of anomalies in dynamic graphs. The black nodes denote the normal nodes, and the red node denotes the abnormality, respectively. The anomaly in the graph becomes noticeable only when analyzing its temporal structural changes rather than just its spatial information from the $t$ snapshot.
@@
Due to the challenge of annotating anomalous objects in real-world scenarios, anomaly detection approaches for dynamic graphs mostly employ an unsupervised learning framework [10]. The key intuition behind these methods is to develop a model that captures patterns of normality by exclusively incorporating normal data during the training phase. Subsequently, objects that the model fails to accurately represent in the testing phase are classified as anomalies. For example, AddGraph [2] employs stacked GCN and GRU to capture spatial and temporal representations of normal edges and train an edge anomaly detector with link prediction as the proxy task. Then the edges with higher prediction errors in the test set are considered abnormal. TADDY [11] encodes representation from dynamic graphs via a dynamic graph transformer in the training phase. The anomalous edges are detected based on a link prediction proxy task similar to [2] in the test phase. NetWalk [6] adopts a random-walk-based encoder to learn node representations and measures the node anomaly score by concerning its closest distance to the normal cluster centers. MTHL [12] projects multi-view dynamic graphs into a shared latent subspace and learns a compact hypersphere surrounding normal samples. Node anomalies are detected based on the distance to the learned hypersphere center. OCAN [13] seizes the normal activity patterns of observed benign users’ attributes and detects fraudsters that behave significantly differently.
@@
Despite the success of these approaches, they encounter two significant limitations: (1) Approaches such as AddGraph[2], TADDY [11], StrGNN[14], etc. only leverage proxy tasks (e.g., edge stream prediction) to derive general feature representations rather than directly identifying normal patterns. These approaches may not be effective if anomalies share commonalities, such as local structures or attributes, with normal data, especially considering the strong expressive power of representation learning models (e.g., GCN, Transformer) used in these studies. Consequently, abnormal data might also be adequately represented, resulting in suboptimal performance for anomaly detection (P1). (2) While methods like NetWalk [6], MTHL [12], OCAN [13], etc. explicitly model normal patterns through clustering or defining hypersphere centers, they do not distinguish the difference between spatial and temporal patterns, but instead treat them as uniform entities. However, as illustrated in Fig. 1, normal data often display distinct prototypical patterns in spatial and temporal dimensions, necessitating separate identification and storage of spatial and temporal normal patterns, respectively (P2).
@@
In the training phase, spatial and temporal memory items are updated together with the encoder and decoder. We propose a comprehensive training objective that not only minimizes reconstruction errors but also reduces compactness errors which promote proximity between node embedding and its nearest memory prototype, and separateness errors which enhance the discriminative power of node embeddings to different memory items. This ensures STRIPE’s effective and judicious use of the limited number of memory items, enabling the accurate extraction of normality prototypes within the training data. In the testing phase, the learned memory items remain fixed, and the comprehensive objective now serves as the anomaly score. Since the reconstruction process integrates normality patterns preserved in memory, inputs that deviate from these prototypical patterns of normal data are likely to yield elevated anomaly scores, thereby facilitating their identification as anomalies.
@@
  • We propose a novel spatial-temporal memories enhanced graph autoencoder framework, STRIPE, that explicitly captures normality patterns and integrates them into graph stream reconstruction for anomaly detection. To the best of our knowledge, this is the first work to integrate memory networks for anomaly detection in dynamic graphs.
@@
  • Extensive experiments have been conducted on multiple benchmark datasets, and the results of the experiment demonstrate that STRIPE achieves state-of-the-art performance with AUCs of $96.20\%$, $97.65\%$, $98.10\%$ and $93.89\%$ on benchmark datasets DBLP-3, DBLP-5, Reddit and Brain, respectively.
@@ II Related Work
In this section, we introduce the works closely related to ours: Anomaly Detection in Dynamic Graphs and Memory Networks.
II-A Anomaly Detection in Dynamic Graphs
@@
More recently, another branch of methods employs deep learning techniques to capture anomalous objects in dynamic graphs [2, 6, 14, 11]. NetWalk [6] utilizes a random walk-based encoder to generate node embeddings and score the abnormality of nodes and edges with their distance to cluster centers. AddGraph [2] employs stacked GCN and GRU to capture spatial and temporal representations of normal edges and train an edge anomaly detector with edge stream prediction as the proxy task. StrGNN [14] further extracts the h-hop enclosing subgraph for each edge and employs stacked GCN and GRU to encode the extracted subgraphs for edge stream prediction. TADDY [11] learns the representations from dynamic graphs with coupled spatial-temporal patterns via a dynamic graph transformer for edge anomaly detection.
@@ II-B Memory Networks
TABLE I: Summary of Notations.
Notations Descriptions
$\mathbb{G}=\{\mathcal{G}^{t}\}^{T}_{t=1}$ A graph stream with maximum timestamp of $T$.
$\mathbf{G}^{t}=\{\mathcal{V}^{t},\mathcal{E}^{t}\}$ The graph snapshot at timestamp $t$.
$\mathcal{V}^{t}$ The node set at timestamp $t$.
$\mathcal{E}^{t}$ The edge set at timestamp $t$.
$v^{t}_{i}\in\mathcal{V}^{t}$ A node with index $i$ at timestamp $t$.
$e^{t}_{i,j}=(v^{t}_{i},v^{t}_{j})\in\mathcal{E}^{t}$ An edge between $v^{t}_{i}$ and $v^{t}_{j}$ at timestamp $t$.
$N^{t}=|\mathcal{V}^{t}|$ Node number at timestamp $t$.
$M^{t}=|\mathcal{E}^{t}|$ Edge number at timestamp $t$.
$\mathbf{A}^{t}$ The adjacency matrix at timestamp $t$.
$\mathbf{X}^{t}$ The node feature matrix at timestamp $t$.
$D$ The dimension of input node features.
$D^{\prime}$ The hidden dimension of node features.
$f(\cdot)$ anomaly detection function.
$\{\mathcal{G}^{t-\tau+1},\ldots,\mathcal{G}^{t}\}$ The sequence of graphs between
timestamps $t-\tau+1$ and $t$.
$\tau$ Window size.
$K_{t}$ The kernel width of gated temporal convolution.
$\alpha$ The balance factor for attributive and structural errors.
$\mathbf{H}^{t}$ The encoded spatial node embeddings at timestamp $t$.
$\mathbf{Z}$ The encoded temporal node embeddings
for $\{\mathcal{G}^{t-\tau+1},\ldots,\mathcal{G}^{t}\}$.
$\mathbf{M}_{sp}$ spatial memory banks.
$\mathbf{M}_{tp}$ temporal memory banks.
$\mathbf{k}_{p}$ The key vector for memory item $\mathbf{m}_{p}$.
$\mathbf{q}_{i}$ The query vector for node $i$.
$\mathbf{v}_{i}$ The value vector for node $i$.
$\hat{\mathbf{M}}_{sp}$ The averaged spatial memory matrix.
$\hat{\mathbf{M}}_{tp}$ The averaged temporal memory matrix.
$\mathbf{Z}$ The decoded temporal node embeddings
for $\{\mathcal{G}^{t-\tau+1},\ldots,\mathcal{G}^{t}\}$.
$\hat{\mathbf{H}}^{(t)}$ The decoded spatial node embeddings at timestamp $t$.
$\mathbf{\Theta}^{(l)}_{en}$ The weight matrix of the l-th layer GNN in encoder.
$\mathbf{\Theta}^{(l)}_{de}$ The weight matrix of the l-th layer GNN in decoder.
$\mathbf{\Phi}^{(l)}_{en}$ The weight matrix of the temporal convolution kernel.
$\mathbf{W}_{K}$ The weight matrix for key vector.
$\mathbf{W}_{Q}$ The weight matrix for query vector.
$\mathbf{W}_{V}$ The weight matrix for value vector.
@@ III Preliminaries
In this section, we provide the definitions of essential concepts and formalize the problem of dynamic graph anomaly detection. We also summarize the frequently used notations across this paper in Table I for quick reference.
Definition 1: Dynamic Graph. Given a dynamic graph with overall timestamps of T, we use $\mathbb{G}=\{\mathcal{G}^{t}\}^{T}_{t=1}$ to denote the graph stream, where each $\mathcal{G}^{t}=\{\mathcal{V}^{t},\mathcal{E}^{t}\}$ is the snapshot at timestamp $t$. $\mathcal{V}^{t}$ and $\mathcal{E}^{t}$ is the node set and edge set at timestamp $t$. An edge $e^{t}_{i,j}=(v^{t}_{i},v^{t}_{j})\in\mathcal{E}^{t}$ indicates the connection between node $v^{t}_{i}$ and $v^{t}_{j}$ at timestamp $t$, where $v^{t}_{i},v^{t}_{j}\in\mathcal{V}^{t}$. $N^{t}=|\mathcal{V}^{t}|$ and $M^{t}=|\mathcal{E}^{t}|$ indicate the number of nodes and edges in timestamp $t$. The structural information of $\mathcal{G}^{t}$ is represented by the graph adjacency matrix $\mathbf{A}^{t}\in\mathbb{R}^{N^{t}\times N^{t}}$, where $\mathbf{A}^{T}_{ij}=1$ if $e^{t}_{ij}$ exists, otherwise $\mathbf{A}^{t}_{ij}=0$. $\mathbf{X}^{t}\in\mathbb{R}^{N^{t}\times D}$ denotes the node feature matrix at timestamp $t$ and its $i$-th row vector $\mathbf{x}^{t}_{i}\in\mathbb{R}^{D}$ represents the feature of node $v^{t}_{i}$.
Definition 2: Anomaly Detection in Dynamic Graphs. Given a dynamic graph $\mathbb{G}=\{\mathcal{G}^{t}\}^{T}_{t=1}$, our goal is to lean an anomaly detection function $f(\cdot):\mathbb{R}^{N^{t}\times D}\rightarrow\mathbb{R}^{N^{t}\times 1}$ that can measure the degree of abnormality of the node by calculating its anomaly score $f(v^{t}_{i,j})$. A larger anomaly score indicates a higher abnormal probability for $v^{t}_{i,j}$.
Given the challenges in obtaining and accessing anomalous labels in real-world networks, we adopt an unsupervised setting for detecting anomalies in dynamic graphs. In the training phase of this research, no node labels that indicate abnormalities are used. Instead, it is presumed that all nodes present during training exhibit normal behavior. The binary labels denoting abnormality are introduced only in the testing phase to assess the model’s effectiveness. In this context, a label $y_{v^{t}_{i,j}}=0$ signifies that the node $v^{t}_{i,j}$ is considered normal, whereas $y_{v^{t}_{i,j}}=1$ identifies it as abnormal.
@@ IV Methodology
In the rest of this section, we introduce the four components in detail from section IV-A to V-E. The overall pipeline of STRIPE is illustrated in Algorithm (1).
@@ IV-A1 Spatial Encoder
[…]
@@
[…]

### v2

# Detecting Anomalies in Dynamic Graphs via Memory enhanced Normality

@@ Abstract
Anomaly detection in dynamic graphs presents a significant challenge due to the temporal evolution of graph structures and attributes. The conventional approaches that tackle this problem typically employ an unsupervised learning framework, capturing normality patterns with exclusive normal data during training and identifying deviations as anomalies during testing. However, these methods face critical drawbacks: they either only depend on proxy tasks for representation without directly pinpointing normal patterns, or they neglect to differentiate between spatial and temporal normality patterns. More recent methods that use contrastive learning with negative sampling also face high computational costs, limiting their scalability to large graphs. To address these challenges, we introduce a novel Spatial-Temporal memories-enhanced graph autoencoder (STRIPE). Initially, STRIPE employs Graph Neural Networks (GNNs) and gated temporal convolution layers to extract spatial and temporal features. Then STRIPE incorporates separate spatial and temporal memory networks to capture and store prototypes of normal patterns, respectively. These stored patterns are retrieved and integrated with encoded graph embeddings through a mutual attention mechanism. Finally, the integrated features are fed into the decoder to reconstruct the graph streams which serve as the proxy task for anomaly detection. This comprehensive approach not only minimizes reconstruction errors but also emphasizes the compactness and distinctiveness of the embeddings w.r.t. the nearest memory prototypes. Extensive experiments on six benchmark datasets demonstrate the effectiveness and efficiency of STRIPE, where STRIPE significantly outperforms existing methods with 5.8% improvement in AUC scores and 4.62$\times$ faster in training time.
@@
Anomaly Detection, Dynamic Graphs, Memory Networks, Graph Autoencoder.
11footnotetext: Corresponding authors.
@@ I Introduction
Real-world networks are often modeled as dynamic graphs to capture the changing nature of objects and their interactions [1, 2, 3, 4]. Beyond basic topology structure and node attributes, dynamic graphs encompass rich temporal signals, such as evolving patterns in graph structure and node attributes [5, 6, 7]. This temporal dimension provides an additional perspective for analyzing anomalies. For instance, as illustrated in Fig. 1, anomalies may not be apparent when considering only the spatial information in a single snapshot, due to the similar degrees of fraudsters and normal users. However, observing temporal changes in graph structures can make abnormalities with higher vibration frequencies distinctly noticeable.
@@
To avoid ambiguity, in this paper, we define the structures and attributive features within individual graph snapshots as spatial information, while characterizing the evolving changes and trends among different snapshots as temporal information. To model and integrate both the spatial and temporal signals of nodes and edges for anomaly detection, there has been a growing interest in the study of anomaly detection in dynamic graphs [8, 9].
Figure 1: Statistical observations of dynamic graphs on DGraph dataset. Left: Average degrees of fraudsters and benign users on snapshot #0. Right: Degree curves of fraudsters and benign users as time evolves. Anomalous samples typically exhibit a higher frequency of vibrations, indicating that fraudsters frequently change their connections to other nodes.
@@
Due to the challenge of annotating anomalous objects in real-world scenarios, anomaly detection approaches for dynamic graphs mostly employ an unsupervised learning framework [10]. The key intuition behind these methods is to develop a model that captures patterns of normality by exclusively incorporating normal data during the training phase. Subsequently, objects that the model fails to accurately represent in the testing phase are classified as anomalies. For example, AddGraph [2] employs stacked GCN and GRU to capture spatial and temporal representations of normal edges and train an edge anomaly detector with link prediction as the proxy task. Then the edges with higher prediction errors in the test set are considered abnormal. TADDY [11] uses a graph transformer to encode the representation of dynamic graphs in the training phase. The anomalous edges are detected based on a link prediction proxy task similar to [2] in the test phase. NetWalk [6] adopts a random-walk-based encoder to learn node representations and measures the node anomaly score by concerning its closest distance to the normal cluster centers. MTHL [12] projects multi-view dynamic graphs into a shared latent subspace and learns a compact hypersphere surrounding normal samples. Node anomalies are detected based on the distance to the learned hypersphere center. OCAN [13] captures the normal activity patterns of observed benign users’ attributes and detects fraudsters that behave significantly differently.
@@
Despite their success, these methods face several limitations: (1) Methods like AddGraph[2], TADDY [11], StrGNN[14] only leverage proxy tasks (e.g., edge stream prediction) to derive feature representation, rather than directly identifying normal patterns. This can be ineffective if anomalies share structural or attributive similarities with normal data, due to the powerful representation learning models (e.g., GCN, Transformer) used. Consequently, abnormal data might also be well-represented, resulting in suboptimal anomaly detection (P1). (2) Approaches like NetWalk [6], MTHL [12], SAD [15] explicitly model normal patterns through clustering or hypersphere centers but do not distinguish between spatial and temporal patterns. However, as illustrated in Fig. 1, the prototypical pattern of nodes within a single snapshot can be different from the pattern of evolving trends of nodes among different snapshots, necessitating separate identification and storage of spatial and temporal normal patterns (P2). (3) State-of-the-art methods such as SAD [15] and CLDG [16] employ contrastive learning to build self-supervised frameworks, relying on extensive negative sampling to prevent model collapsing. However, the generation and processing of a substantial number of negative samples in large-scale graphs will bring high computational costs and reduce training efficiency (P3).
@@
During training, spatial and temporal memory items are updated along with the encoder and decoder. We propose a comprehensive training objective that minimizes both reconstruction errors and compactness errors, promoting proximity between node embeddings and their nearest memory items. Additionally, we also minimize separateness errors to enhance the distinctiveness of memory items. This ensures STRIPE’s effective use of a limited number of memory items, significantly reducing the size of the memory bank and enhancing model efficiency. Moreover, since STRIPE is non-contrastive learning model, it requires no negative sampling or data augmentation, which also helps promote model scalability.
In the testing phase, the learned memory items remain fixed, and the reconstruction and compactness loss now serve as the anomaly score. Since the reconstruction process integrates normality patterns preserved in memory, the inputs that deviate from these prototypical patterns of normal data are likely to yield elevated anomaly scores, thereby facilitating their identification as anomalies.
@@
  • We propose a novel spatial-temporal memory-enhanced graph autoencoder framework, STRIPE, that explicitly captures normality patterns and integrates them into graph stream reconstruction for anomaly detection.
@@
  • We propose an efficient dynamic anomaly detection model by promoting separateness among memory items, thereby significantly reducing the size of memory bank.
  • Extensive experiments on six benchmark datasets have demonstrated the state-of-the-art performance of STRIPE, achieving an average AUC score improvement of 5.8%. Both theoretical analysis and empirical results highlight STRIPE’s efficiency, showing linear scalability with the increase of node numbers.
@@ II Related Work
In this section, we introduce the works closely related to ours: Anomaly Detection in static graphs, Anomaly Detection in Dynamic Graphs and Memory Networks.
II-A Anomaly Detection in static graphs
Graph anomaly detection aims at identifying anomalous graph objects (i.e., nodes, edges, and subgraphs) in the graph [8, 9, 18]. Early anomalous node detection approaches mainly use shallow techniques such as residual analysis (Radar [19]), matrix factorization (ALAD [20]), and CUR decomposition (ANOMALOUS [21]) to extract anomalous pattern in graphs. More recently, DOMINANT [22] pioneered the integration of deep learning into node anomaly detection by employing a graph autoencoder [23] to reconstruct both the structure and attribute information of graphs. CoLA [11] SL-GAD [24] and CONAD [25] further introduce graph contrastive learning that captures abnormal patterns by measuring agreement between augmented item pairs.
II-B Anomaly Detection in Dynamic Graphs
@@
More recently, another branch of methods employs deep learning techniques to capture anomalous objects in dynamic graphs [2, 6, 14, 11]. NetWalk [6] utilizes a random walk-based encoder to generate node embeddings and score the abnormality of nodes and edges with their distance to cluster centers. AddGraph [2] employs stacked GCN and GRU to capture spatial and temporal representations of normal edges and train an edge anomaly detector with edge stream prediction as the proxy task. StrGNN [14] further extracts the h-hop enclosing subgraph for each edge and employs stacked GCN and GRU to encode the extracted subgraphs for edge stream prediction. TADDY [34] learns the representations from dynamic graphs with coupled spatial-temporal patterns via a transformer. SAD [15] and CLDG [16] introduce contrastive learning for anomaly detection in dynamic graphs.
@@ III Preliminaries
In this section, we provide the definitions of essential concepts and formalize the problem of dynamic graph anomaly detection.
Definition 1: Dynamic Graph. Given a dynamic graph with overall timestamps of T, we use $\mathbb{G}=\{\mathcal{G}^{t}\}^{T}_{t=1}$ to denote the graph stream, where each $\mathcal{G}^{t}=\{\mathcal{V}^{t},\mathcal{E}^{t}\}$ is the snapshot at timestamp $t$. $\mathcal{V}^{t}$ and $\mathcal{E}^{t}$ is the node set and edge set at timestamp $t$. An edge $e^{t}_{i,j}=(v^{t}_{i},v^{t}_{j})\in\mathcal{E}^{t}$ indicates the connection between node $v^{t}_{i}$ and $v^{t}_{j}$ at timestamp $t$, where $v^{t}_{i},v^{t}_{j}\in\mathcal{V}^{t}$. $N_{t}=|\mathcal{V}^{t}|$ and $M_{t}=|\mathcal{E}^{t}|$ indicate the number of nodes and edges in timestamp $t$. The structural information of $\mathcal{G}^{t}$ is represented by the graph adjacency matrix $\mathbf{A}^{t}\in\mathbb{R}^{N_{t}\times N_{t}}$, where $\mathbf{A}^{t}_{ij}=1$ if $e^{t}_{ij}$ exists, otherwise $\mathbf{A}^{t}_{ij}=0$. $\mathbf{X}^{t}\in\mathbb{R}^{N_{t}\times D}$ denotes the node feature matrix at timestamp $t$ and its $i$-th row vector $\mathbf{x}^{t}_{i}\in\mathbb{R}^{D}$ represents the feature of node $v^{t}_{i}$.
Definition 2: Anomaly Detection in Dynamic Graphs. Given a dynamic graph $\mathbb{G}=\{\mathcal{G}^{t}\}^{T}_{t=1}$, our goal is to lean an anomaly detection function $f(\cdot):\mathbb{R}^{N_{t}\times D}\rightarrow\mathbb{R}^{N_{t}\times 1}$ that can measure the degree of abnormality of the node by calculating its anomaly score $f(v^{t}_{i})$. A larger anomaly score indicates a higher abnormal probability for $v^{t}_{i}$.
Given the challenges in obtaining and accessing anomalous labels in real-world networks, we adopt an unsupervised setting for detecting anomalies in dynamic graphs. In the training phase of this research, no node labels that indicate abnormalities are used. Instead, it is presumed that all nodes present during training exhibit normal behavior. The binary labels denoting abnormality are introduced only in the testing phase to assess the model’s effectiveness. In this context, a label $y_{v^{t}_{i}}=0$ signifies that the node $v^{t}_{i}$ is considered normal, whereas $y_{v^{t}_{i}}=1$ identifies it as abnormal.
@@ IV Methodology
In the rest of this section, we introduce the four components in detail from section IV-A to V-D. The overall pipeline of STRIPE is illustrated in Algorithm (1).
@@ IV-A1 Spatial Encoder
$\mathbf{H}^{(t,l)}=\textit{GNN}\Bigl{(}\mathbf{A}^{t},\mathbf{H}^{(t,l-1)};\mathbf{\Theta}_{en}^{(l)}\Bigr{)},$ (1)
where $\mathbf{\Theta}_{en}^{(l)}\in\mathbb{R}^{D\times D^{\prime}}$ denotes the learnable weight parameters of the $l$-th layer GNN. $\mathbf{H}^{(t,l-1)}$ and $\mathbf{H}^{(t,l)}$ are the node representation matrices learned by the $(l-1)$-th and $(l)$-th layer, respectively. $\mathbf{H}^{(t,0)}$ is $\mathbf{X}^{t}$. $\textit{GNN}_{\theta}(\cdot)$ can be set as any off-the-shelf graph neural networks. For computation efficiency, we adopt a two-layer graph convolutional network (GCN) as the backbone. Thus, Equation (1) can be specifically re-written as:
@@
where $\widetilde{\mathbf{A}}^{t}=\mathbf{A}^{t}+\mathbf{I}_{N^{t}}$, and $\widetilde{\mathbf{D}}_{t}$ is a diagonal node degree matrix where $\widetilde{\mathbf{D}}_{t}(i,i)=\sum_{j}\widetilde{\mathbf{A}}_{t}(i,j)$. $\textit{ReLU}(\cdot)$ is the activation function. We simplify $\mathbf{H}^{(t,L)}$ as $\mathbf{H}^{t}$.